OpenStudy (anonymous):

Using complete sentences, explain each step in simplifying the ratio x squared plus 8 times x all over x squared plus 10 times x plus 16. and provide the simplified form.

OpenStudy (anonymous):

I need help because I literally have no clue how to solve this.

OpenStudy (anonymous):

I believe this is the correct equation. \[\frac{x^{2} + 8x}{x^{2}+10x+16}\] This can be simplified to: \[\frac{x(x + 8)}{(x+8)(x+2)}\] and then we cancel out like things: \[\frac{x}{(x+2)}\] and this is our simplified form. I'll let you come up with the sentences to explain this but just remember we factored the bottom.